Product

2004

2005

2006

Quantity

Price

Quantity

Price

Quantity

Price

Maize

500

20

510

25

550

27

Cotton

220

10

600

12

580

18

Cassava

800

30

700

35

750

35

 

                 ii.           Calculate real GDP for each year using 2004 as the base year.

               iii.           Calculate the GDP deflator for each year taking 2004 as the base year.

               iv.           Calculate the GDP growth rates for the years 2005 and 2006.


1
Expert's answer
2021-07-08T15:15:11-0400

i. real GDP for each year using 2004 as the base year.

Real GDP = Quantity "\times" price of the base year

"2004 \\ (500\times 20) + (220\times 10) + (800\times 30) = \$36200\\

2005 \\ (510\times 20) + (600\times 10) + (700\times 30) = \$37200\\

2006 \\(550\times 20) + (580\times 10) + (750\times 30) = \$39300"


ii. Calculate the GDP deflator for each year taking 2004 as the base year.

GDP deflator = [Nominal GDP "\div" Real GDP (base year)] "\times" 100

"2004 \\ (\frac{36200 } {36200}) \times 100 = 100\\

2005 \\ (\frac{44450 } {36200}) \times 100 = 123\\

2006 \\ (\frac{51540 }{36200}) \times 100 = 142"

Calculation of Nominal GDP = Quantity "\times" Price

"2004 \\(500\times 20) + (220\times 10) + (800\times 30) = \$36200\\

2005 \\ (510\times 25) + (600\times 12) + (700\times 35) = \$44450\\

2006 \\(550\times 27) + (580\times 18) + (750\times 35) = \$51540"


iii. Calculate the GDP growth rates for the year 2005 and 2006.

GDP growth rate = (GDP of present year - GDP of base year} "\div" GDP of base year

"2005 \\\frac{\$37200 - \$36200} { \$36200} = 2.76\%\\

2006 \\\frac{ \$39300 - \$36200 }{ \$36200} = 8.56\%"